Find the equation of a quadratic polynomial passing through the following points: (-1,2) (-5,2) and (6,8)?
Answer by ankor@dixie-net.com(22740)

Quadratic_Equations/52883 (2006-09-19 14:53:26):
:
Find the equation of a quadratic polynomial passing through the following points: (-1,2) (-5,2) and (6,8)?
Using the form: ax^2 + bx + c = y, write an equation for each x,y coordinates
and find the values of a, b, c
There is an equation that closely satisfies the problem, however a,b,c are not
integers. Could there be a type here??
:
a(-1^2) + b(-1) + c = 2
a - b + c = 2
:
a(-5^2) + b(-5) + c = 2
25a - 5b + c = 2
:
a(6^2) + b(6) + c = 8
36a + 6b + c = 8
:
:
Using a matrix to solve:
+1 +1 +1 +2
25 -5 +1 +2
36 +6 +1 +8

Using the matrix feature on a TI83 we get: y = .078x^2 + .467x + 2.39